OpenStudy (anonymous):

simplify the radical: radical 50

OpenStudy (anonymous):

\[5\sqrt{2}\]@babeekatj Are you understanding these?

OpenStudy (anonymous):

not really only one i really understood is radical 144

OpenStudy (campbell_st):

\[\sqrt{50} = \sqrt{2 \times 25} = \sqrt{2} \times \sqrt{25} = 5\sqrt{2}\]